How long must a train undergo constant acceleration in order to go from 35 km/h east to 45 km/h west ending up 250 mm west of its original position ?


1
Expert's answer
2022-02-18T08:22:13-0500

The displacement of a train

"d=\frac{v_i+v_f}{2}t"

The time

"t=\frac{2d}{v_1+v_f}=\frac{2*0.25\rm m}{(-35+45)/3.6\:\rm m/s}=0.18\:\rm s"